April 22, 2010 FDC meeting
- Updates on:
- Electronics: PCB, cable procurement status (Fernando)
- Cathode redesign (Roger)
- FDC cabling (Bill)
- Full-scale prototype
- status (Lubomir)
- grounding problems (Fernando)
- cosmics tests (Beni)
Participants: Bill, Roger, Brian, Glenn, Chuck, Mark, Simon, Beni, Eugene, Lubomir
- Brian reported on his visit at UVA. The parts at UVA, two pin rails (~2m each) and the parts of the position measurement system, are in good condition and can be used. Brian thinks that the stringing of the FDC at JLab using the UVA procedure/equipment is doable; the number of the wires is not that big to require winding machine. Brian and Bill discussed some possible improvements related to the way the pin rails are attached to the table/ strongback, the way the weights are attached to the wires. There were discussions about the precision of the stepper motor for the position measurements.
- not only UVA agreed to lend us the equipment, but as Prof. Nilanga Liyanage (before UVE he worked for Hall A) proposed, they want to join the Hall D collaboration and work on the FDC (great news!). As part of their contribution, Nilanga proposed to send us a student for this summer. We are discussing now the possible tasks for him.
- Glenn suggested to organize a meeting with Rusty to formulate our requirements about the off-site space. Bill already made a floor plan and estimated some general parameters that were needed to plan the infrastructure there.
- Electronics: the PCBs are finalized; Kim sent out the PCB specifications to three vendors. Fernando submitted the PR for the cables.
- Cathode redesign: Roger had several questions related to the ground of the cathode daughter board (flex) design, and also the strip labels.
- Bill explained the proposed design of the cable bundles and the installation procedure. Still, we have to decide about the total number of cables and get accurate info about their sizes. This includes 528 (+ 24/48spare?) signal cables, 264 thin low voltage cables, 8 HV cables, 4 grounding cables, 48 gas pipes, 8? cooling lines. We are considering also to do some prototype cable bundles to make sure it works. Expect more discussion on the next meeting.
Full scale prototype
- At the end of the last week, the the three-layer package was lifted from the table, holding it only on the two supporting brackets to check for gas leakages due to possible deformations. After it passed this test, we mounted the chamber at the testing set-up. We were able to bubble the three layers again at 210cc/min. We had several runs with comics.
- We have a problem: there are oscillations at the output of the pre-amps especially at the cathode cards. The oscillations are stronger in amplitude for the longest strips. Fernando, Beni and Lubomir have tried different configurations to investigate the problem: old/new cards, different kinds of shielding both of the detector and of the cards, improving the ground on the cards, etc. For now, there are no conclusive results. Fernando found that there's an external noise with the same frequency that can be picked up by the strips. The origin of this noise has not been identified. We will make use of the power outage on Friday, to check if the oscillations are there when the power is down; we will use the UPS for the low voltage and the scope. We agreed, the study of the oscillations will be our priority. If necessary, on Monday we will move the chamber back in the tent, where it is easier to investigate the problem.
- Beni did several tests with cosmics at the top and middle layers. Only the information from the wires is useful because of the above problem. He's going to correlate the tracks form the IU chamber and the hits at our chambers, and also to look at the wire response of the top two chambers that we installed with the wires parallel to each other.
- Eugene wanted to have the latest floor plan and main parameters for the off-site space. He also wanted to make sure we have a task for the UVA student this summer.